OpenStudy (anonymous):

You pay 10% of the total price of the stock and borrow the rest. Stock in American Airlines cost $200 a share, I have $2000 dollars but I want 100 shares. a) How much do I pay? b) How much do I borrow? c) What's the total? d) Number of shares?

OpenStudy (triciaal):

100 shares @ $200 each means cost is 100x 200 = $20000 pay 10% of 20000 = (10/100)(20000) = $2000 borrow (100%-10%)= 90% 90% of 20000 = (90/100)(20000) (same as 20000-2000 the total cost - what you pay) so you borrow $18000 $2000/$200 per share = 10 shares